### v2

# Electronic origin of reorganization energy in interfacial electron transfer

Electronic origin of reorganization energy in interfacial electron transfer
Electron transfer (ET) reactions underpin energy conversion and chemical transformations in both biological[1, 2, 3] and abiological[4, 5, 6, 7] systems. The efficiency of any ET process relies on achieving a desired ET rate within an optimal driving force range. Marcus theory[8, 9] provides a microscopic framework for understanding the activation free energy, and thus the rate, of ET in terms of a key parameter: the reorganization energy. For electrified solid–liquid interfaces, it has long been conventionally understood that only factors in the electrolyte phase are responsible for determining the reorganization energy and the electronic density of states (DOS) of the electrode serves only to dictate the number of thermally accessible channels for ET.[10, 7, 11, 12, 13, 14] Here we show instead that the electrode DOS plays a central role in governing the reorganization energy, far outweighing its conventionally assumed role. Using atomically layered heterostructures, we tune the DOS of graphene and measure outer-sphere ET kinetics. We find the ensuing variation in ET rate arises from strong modulation in a reorganization energy associated with image potential localization in the electrode. This work redefines the traditional paradigm of heterogeneous ET kinetics, revealing a deeper role of the electrode electronic structure in interfacial reactivity.
In the initial formulation of Marcus theory for homogeneous ET involving redox-active ions in solution, the activation free energy was explained in terms of a reorganization energy () penalty required to distort the atomic configuration and solvation environment of the reactant species to resemble those of the product state[8, 9]. Extensions in the so-called Marcus–Gerischer[10] and Marcus–Hush[7, 11] formalisms rationalized heterogeneous ET processes at electrode–electrolyte interfaces, specifically addressing the ET rate constant in the weak coupling limit. The quantum mechanical theory of ET, pioneered by Levich, Dogonadze, Chernenko, and Kuznetsov [13, 14], similarly leads to Marcus–Hush-type rate expressions. The seminal adaptation of the Marcus–Hush model by Chidsey[12] explained the dependence of interfacial ET rates on driving force and temperature by incorporating the Fermi–Dirac distribution of occupied electronic states in the electrode. In all these and later[15, 16] extensions that incorporated a non-uniform (energy-dependent) density of electronic states (DOS) profile of the electrode, the DOS of the electrode serves exclusively to dictate the number of thermally accessible channels for ET. Additionally, consistent with the original framework, is presumed to arise largely from nuclear reconfigurations in the electrolyte phase (typically those of the solvent and in some cases the redox molecule itself).
Enhancements in interfacial charge transfer at electrodes/photoelectrodes due to electrostatic variations in carrier doping or defects (including vacancies, step edges, and grain boundaries) are commonly explained as arising from increases in the electrode DOS near the Fermi level (), ostensibly due to increased number of thermally accessible channels for ET[17, 18, 19]. Yet the Marcus–Hush–Chidsey/Marcus–Gerischer framework often fails to quantitatively predict interfacial ET rate constants, even for relatively simple electrode reactions, overestimating ET rates by an order of magnitude or more[20, 21, 22, 23]. A recent example of these discrepancies is found in the interfacial ET behavior of twisted bilayer graphene and twisted trilayer graphene[16, 24]. In these moiré electrode systems, which display periodic spatial localization of the electronic charge density in moiré superlattice topological defects, models derived from the Marcus–Hush–Chidsey (MHC) framework are unable to account for the large variation in ET rate with twist angle despite modifications to account for a DOS profile that varies with energy[15] as well as quantum capacitance effects that lead to changes in DOS of the electrode upon electrochemical polarization[16, 24].
What is missing from these frameworks to explain how electrode defects produce such strong local enhancements in interfacial ET is a consideration of the electrode’s contribution to . Yet, molecular dynamics simulations have long predicted that for interfacial ET would vary with the distance of a redox molecular ion from a metallic electrode, owing to image charge interactions[21, 25, 26]. First-principles calculations of ET rates at graphite electrodes also considered a contribution of the electrode dielectric response to , but the effect was presumed to be sufficiently small to be neglected[27]. Moreover, recent simulations of twisted bilayer graphene have established a connection between the moiré twist angle and the screening of charge carriers within the electrode, identifying that a twist-angle dependent reorganization energy can account for the interfacial ET behavior observed at moiré graphene electrodes[28, 29]. A unique aspect of these electrode systems is the ability to continuously change the density of states at the Fermi energy, and correspondingly tune the ability of the electrode to screen electric fields. The Thomas-Fermi screening length, , quantifies the lengthscale over which charges are screened in imperfect metals. Because scales inversely with DOS, higher metallicity leads to sharper charge localization, whereas lower metallicity yields a more diffuse charge distribution. Such tunability offers a new avenue to investigate how electronic screening shapes interfacial ET.[30]
In this study, we directly and systematically probe the DOS dependence of interfacial ET using van der Waals (vdW) assembly of atomically thin crystals. Employing solid state dopant layers and hexagonal boron nitride (hBN) spacers, we electrostatically tune the doping levels in monolayer graphene and measure the resulting variation in heterogeneous outer-sphere electrochemical ET rates of the [Ru(NH couple. The strong variation in graphene charge density with changes to hBN spacer thickness is shown to be mediated by defects in the hBN crystals. A continuum model is leveraged to obtain a microscopic understanding of the dependence of interfacial ET rate on the electrode DOS. We find that the ensuing variation in ET rates with charge carrier density cannot be modeled in the Marcus framework when one only considers the change in the number of thermally accessible channels for ET. Instead, our measurements and simulations unveil the considerably more dominant DOS-dependent reorganization energy, which accurately captures the large experimental variation in interfacial ET rate with DOS. We observe that at low charge carrier densities, such as those found in many low-dimensional electrode materials as well as bulk or nanocrystalline semiconductors, the reorganization energy penalty owing to the low electrode DOS can be of magnitude comparable to that arising in the solvent at a metallic electrode. This systematic study of the DOS dependence of interfacial ET rates on well-defined electrode surfaces challenges the conventional paradigm that reorganization energy contributions predominantly arise from the electrolyte side of the electrode–electrolyte interface and establishes a general microscopic framework for understanding heterogeneous ET that explicitly accounts for the electronic properties of the electrode in governing the free energy of activation.
Measurements of interfacial ET as a function of DOS
Isolating two-dimensional (2D) crystals enables the assembly of vdW heterostructures with tailored electronic and chemical properties[31, 32, 33, 34]. When 2D crystals with disparate work functions are interfaced, the resulting electric field leads to charge transfer and interfacial doping, analogous to the effect of an electrostatic gate in a field effect transistor, and prior work has shown that -RuCl3 and WSe2 can thus be used to dope graphene with holes and electrons, respectively[35]. This approach provides a modular doping mechanism similar to vacancy or substitutional doping but avoids introducing chemical disorder into the active layer. Such heterostructures provide an exceptionally well-defined platform for examining how doping-induced DOS changes impact the rate of ET.
We fabricated mesoscopic electrochemical devices comprising monolayer graphene (MLG) on RuCl3. We also fabricated samples with hexagonal boron nitride (hBN) spacers placed between MLG and RuCl3, and varied the hBN spacer thickness from 3 nm to 120 nm (see Materials and Methods for additional details on sample fabrication). Electrochemical measurements were conducted using scanning electrochemical cell microscopy (SECCM), which enables nanoscale electrochemical measurements by positioning an electrolyte-filled nanopipette over the sample and forming a confined electrochemical cell upon meniscus contact[36]. In this study, we employed quartz nanopipettes of 600–800 nm diameter (Extended Data Fig. S3) containing 2 mM hexaammineruthenium(III) chloride and 100 mM potassium chloride as supporting electrolyte.
Figures 1a,b present schematics of the sample and measurement setup alongside an optical micrograph of a representative MLG/hBN/RuCl3 device in contact with a graphite flake, which serves as the electrical contact. Steady-state cyclic voltammograms (CVs) of the [Ru(NH3) couple at the basal plane of a MLG/10 nm-hBN/RuCl3 heterostructures (Figure 1c) reveal a shift in the half-wave potential to more positive potentials compared to CVs of analogous samples without RuCl3, consistent with enhanced DOS from RuCl3-induced hole doping that facilitates the electroreduction of [Ru(NH3). Notably, even with a 10 nm (25-layer) hBN spacer, MLG displays ET kinetics approaching those of graphite and exhibits nearly reversible electrochemical behavior at the basal plane. This enhanced kinetic behavior can be understood as originating from a downward shift in the relative to the charge neutrality point (CNP), driven by the work function difference between RuCl3 and MLG. To a first approximation, the associated increased DOS (hole doping) at would substantially expand the availability of states to mediate interfacial ET, increasing the extent of energy overlap between states in Ru(NH3) and MLG, in the Gerisher–Marcus framework.
Finite-element simulations were performed using COMSOL Multiphysics v.5.4 to simulate voltammetric responses with the Butler–Volmer model and estimate standard electrochemical rate constants (), as detailed in the Methods. Figure 1d illustrates the correlation between values and hBN thickness for RuCl3 (solid-state acceptor) and WSe2 (solid-state donor), revealing a strong modulation in ET kinetics with varying hBN thickness. Even when a 50 nm thick hBN spacer layer is used, RuCl3 leads to a measurable increase in relative to the electrochemical response measured in the absence of RuCl3. In the case of WSe2, the effect diminishes beyond an hBN spacer thickness of 20 nm. In the absence of an hBN spacer, MLG/RuCl3 achieves ET rates comparable to those of pristine bulk graphite.
Carrier density and fluorescence measurements
Raman spectroscopy and Hall measurements were used to quantify doping as a function of hBN thickness[35]. Figure 2a (top) compares the Raman G-peak positions of pristine MLG with those of MLG/RuCl3, MLG/WSe2, and MLG/hBN/RuCl3 heterostructures across hBN thicknesses (Figure 2a, bottom). The doping-induced shifts in the graphene G-peak position are stronger for RuCl3 than WSe2 and are seen to attenuate with increasing hBN thickness. Prior work has shown a linear relationship between G peak position and charge carrier density[35], establishing a doping change of about cm-2 carriers per wavenumber shift. This correlation was used to deduce the graphene charge carrier density induced by RuCl3 or WSe2 heterolayers as a function of hBN spacer thickness.
When MLG and RuCl3 are separated by hBN crystals thicker than 20 nm, we are unable to resolve further shifts in the G-peak. Since we observe modulations in ET kinetics beyond 20 nm (Figure 1d), we fabricated mesoscopic Hall bar devices and performed electronic transport measurements to measure these low carrier densities (see Methods). Figure 2b presents the Hall resistance () at 1.8 K for MLG/hBN/RuCl3 devices with hBN thicknesses of 22, 43, and 79 nm. The slope of as a function of magnetic field at low fields is inversely proportional to the carrier concentration, , and indicates carrier type (positive: holes, negative: electrons), allowing for direct assessment of doping levels. Up to an hBN spacer thickness of 43 nm, RuCl3 generates hole doping in MLG, whereas a negative slope is measured in the pristine MLG (no RuCl3) region of the same device as well as in the MLG/hBN/RuCl3 device consisting of 79 nm hBN, indicating electron doping.
Figure 2c plots in MLG as a function of hBN thickness, derived from Raman G-peak shifts (circles) and Hall measurements (squares). These experimental values are compared to the doping expected from first-principles calculations[37] (black dashed line) that consider the work function difference between MLG and RuCl3 and dielectric constants of the spacers. While the model aligns very well with experimental data for hBN thicknesses 20 nm, doping levels exceed those predicted by theory for thinner hBN spacer layers. The enhanced doping observed in ultrathin hBN may be a result of valence band alignment between hBN and RuCl3, which activates a theoretically proposed defect-mediated charge transfer process[38]. To investigate this possibility of hBN defect-mediated doping, we conducted fluorescence measurements on hBN-RuCl3 heterostructures (Figure 2e). Prior work has shown that organic solvents enhance hBN defect fluorescence via solvent-defect charge transfer[39]. Using wide-field imaging (see Methods), we observed that RuCl3 strongly quenches the defect-based fluorescence in hBN (Figure 2f). This result suggests that defect-mediated interactions may indeed contribute to the increased extent of MLG doping by RuCl3, which is observed when hBN 20 nm thick is used in MLG/hBN/RuCl3 heterostructures.
Effect of electrode metallicity on the reorganization energy
By relating the hBN-modulated carrier densities (Figure 2c) to shifts in the chemical potential, , (i.e., at equilibrium) of MLG relative to its band structure, Figure 3a compares experimental outer-sphere standard ET rate constants (symbols) as a function of to theoretical models (lines). In our modeling (see Materials and Methods), rate constants are normalized to those at the undoped graphene (), and are calculated at zero overpotential. Under these conditions, quantum capacitance effects on are negligible compared to the effects of solid-state doping.
First, we consider an MHC model (red dotted line in Figure 3a) in which the reorganization energy (metal) is assumed to be a constant = 0.82 eV, as measured for this redox pair on metal (gold modified by self-assembled monolayer thiols) electrodes[40]. This MHC model accounts for the doping dependent DOS but captures only the effect on interfacial ET of increasing the number of thermally accessible states in the electrode. The result is a prediction of very modest rate enhancement with DOS relative to that at the undoped graphene ( 5–10-fold at extreme doping) that completely fails to replicate the experimental scaling.
Instead, we find that the strongest contribution to the ET rate enhancement with increasing DOS is derived from an attenuation in the reorganization energy. This DOS dependence of arises in response to changes in the dielectric screening within the electrode as a function of doping, and can be understood in terms of image charge interactions at the electrochemical interface. To explain this effect, Figure 3b presents a series of simulation snapshots of the instantaneous discharging of an empty capacitor, displaying logarithmic maps of the induced charge distributions in response to a monovalent ion placed 5 Å away from the surface. The simulations were performed subject to a constant potential constraint, where electron-electron screening was modeled within Thomas-Fermi theory [41, 30, 42]. The screening length, , is inversely proportional to the Fermi level, , reflecting enhanced metallicity as the charge carrier density increases. As increases (corresponding to an increased metallicity and decrease in ), the induced charge density becomes sharply localized, whereas a decrease in doping, or equivalently a decrease in metallicity, leads to a more diffuse charge distribution.
Next, in Figure 3c, we account for this to model a reorganization energy, , that is a function of DOS and using non-local dielectric continuum theory. Our model predicts a strong attenuation of with increasing metallicity (increasing DOS), converging to a value of , consistent with metal[40]. This trend arises from enhanced stabilization of the charge transfer transition state due to sharply localized polarization within the electrode. This manifests in the calculated free energy surfaces for the Ru3+/Ru2+ redox system at different values of (left inset of Figure 3c). These calculations reveal a pronounced influence of electrode screening on electron transfer kinetics: reduced monotonically increases the activation free energy (), corresponding to the intersection of the reactant and product free energy surfaces. The right inset of Figure 3c plots vs. . The plot reveals a rapid decline in as the system is doped away from the CNP (), directly linking doping-induced Fermi level shifts to screening efficiency. By explicitly including —which decreases with rising DOS—our model achieves quantitative agreement with experimental rates (red line in Figure 3a), establishing that electrode electronic structure governs and by extension the activation free energy landscape. Therefore, the predominant contributor to increasing ET rates with increasing DOS is the decrease in , not the change in number of thermally accessible electron donor/acceptor states.
Conclusions
Our experimental and theoretical results demonstrate the pivotal influence of the electrode electronic structure on interfacial ET kinetics, not only through the expected increase in the pre-exponential factor as DOS increases, but also through a significant impact on the reorganization energy itself. This paradigm shift challenges the traditional electrolyte-centric perspective of reorganization dynamics, compelling a revised activation free energy framework that explicitly integrates the influence of electrode electronic properties on . Our findings provide a more comprehensive mechanistic understanding of heterogeneous electron transfer and establish foundational principles to guide the design and optimisation of interfacial charge transfer in next-generation devices. We anticipate that this concept will be particularly relevant to photo-induced ET processes, which inherently involve semiconducting (thus low DOS) materials[43, 44, 45]. Future theoretical and experimental efforts should further unravel the interplay between electrode DOS, defect engineering, and to fully harness the potential of low-DOS materials in quantum technologies and energy applications.
Acknowledgements
This material is based upon work supported by the US Department of Energy, Office of Science, Office of Basic Energy Sciences, under award no. DE-SC0026007 (experimental studies by S.M., Y.W., M.P.E., J.N.N., and D.K.B.) and DEAC02-05-CH11231 within the Fundamentals of Semiconductor Nanowire Program (KCPY23) (theoretical work by L.C.E. and D.T.L.). Confocal Raman spectroscopy was supported by a Defense University Research Instrumentation Program grant through the Office of Naval Research under award no. N00014-20-1-2599 (D.K.B.). Fluorescence measurements by A.T. and K.X. were supported by STROBE, a National Science Foundation Science and Technology Center under Grant No. DMR 1548924. K.W. and T.T. acknowledge support from the JSPS KAKENHI (Grant Numbers 21H05233 and 23H02052) , the CREST (JPMJCR24A5), JST and World Premier International Research Center Initiative (WPI), MEXT, Japan.
Author Contributions
SM and DKB conceived the study. SM and YW fabricated the samples with assistance from MPE. JNN grew the RuCl3 crystals. SM and YW performed the electrochemical measurements and COMSOL simulations. LCE and DTL carried out the theoretical modeling. SM and MPE carried out the transport measurements. SM and AT performed the fluorescence measurements under the supervision of KX. TT and KW provided the hBN crystals. SM and DKB analyzed the data. SM, LCE, DTL, and DKB wrote the paper.
Competing Interests
The authors declare no competing interests.
Additional Information
Correspondence and requests for materials should be emailed to DKB (email: bediako@berkeley.edu) and DTL (dlimmer@berkeley.edu).